The length of a rectangle is 2 feet less than 3 times the width. The sum of the length and width is 13 more than the width. Find the width.
Answer by josgarithmetic(39617)

w for width and L for length;

L=-2+3w; L+w=13+w; those are the equations transcribed from the description, and those are enough to find L and w. Simplifying the second equation and then substituting for L in it is a good (but not the only) way to go.